The perimeter of a rectangular flower garden is 60 m and its area is 225 m^2. Find the length of the garden. Use L = length and W = width.
A. Write a quadratic equation in terms of L in the form 0 = aL^2 + bL + c that represent the area of the flower bed.
B. Solve the quadratic equation you wrote in part a and say which method you used.
Please answer showing all steps i am behind and getting ready to do a test on this and i didn't study for this paticular promblem please help!

Answer by stanbon(75887)

The perimeter of a rectangular flower garden is 60 m and its area is 225 m^2. Find the length of the garden. Use L = length and W = width.
======
A. Write a quadratic equation in terms of L in the form 0 = aL^2 + bL + c that represent the area of the flower bed.
60 = 2L + 2W
225 = L*W
--------------
L + W = 30
L*W = 225
-----
Substitute for "W" to get:
L(30-L) = 225
-L^2 + 30L - 225 = 0
==========================
B. Solve the quadratic equation you wrote in part a and say which method you used.
L^2 - 30L + 225 = 0
Quadratic Equation:
L = [30 +- sqrt(900-4*1*225)]
That gives rise to an imaginary answer; please check your post.
Cheers,
Stan H.
-------------